OpenStudy (anonymous):

Triangles MOP and MNQ are similar. Find x. NQ is 5,OP is x ,MN is 15 and NO is 3

OpenStudy (anonymous):

Equation: NQ / OP = MN / NO Substituting the values: 5 / (5 + x) = 15 / (15 + 3) Cross multiplication: x = ?

OpenStudy (anonymous):

wouldnt it be 5/x=15/3? then i cross multiply?

OpenStudy (anonymous):

Oh my bad. I suppose I put it wrong :o but yes you cross multiply and you should find x.

OpenStudy (anonymous):

thank you, i guess x is 1